# Learning More by Seeing Less: Line Drawing Pretraining for Efficient, Transferable, and Human-Aligned Vision

### v1

@@ Results
To further assess the inductive biases learned by our models, we evaluated shape versus texture preferences using the model-vs-human benchmark (Geirhos et al. 2021). This benchmark presents cue-conflict images in which shape and texture cues are intentionally mismatched, allowing for a direct comparison of classification strategies between models and human observers.
Figure 7 displays the distribution of shape and texture decisions across object categories (these are the 7 categories in the cue-conflict dataset that coincide with STL10 classes). Models pretrained on line drawings and subsequently fine-tuned on color images demonstrate a pronounced shift toward shape-based decisions, as indicated by the clustering of yellow circles toward higher fractions of shape choices. In contrast, models trained only on natural images show a greater reliance on texture cues, with more blue circles distributed toward the texture-dominant end of the spectrum. This pattern suggests that line drawing pretraining encourages the model to prioritize global structural features over local textural information when resolving ambiguous visual input.
These results indicate that line drawing pretraining not only enhances classification performance and attention focus, but also aligns model decision-making more closely with human perceptual strategies. This increased shape bias likely contributes to the improved robustness and generalization observed in downstream tasks, as models become less reliant on superficial texture cues and more attuned to the underlying structure of visual objects.
Benefits in Semantic Segmentation
We further show that pretraining on line drawings enhances transferability to dense prediction tasks such as semantic segmentation. To evaluate this, we fine-tune ResNet-18 models—initialized either from color-only training or from line-drawing → color training—on the ADE20K benchmark.
As shown in Figure 8, the model pretrained with line drawings consistently outperforms the color-only baseline across multiple semantic categories. The improvement is especially pronounced for object-centric classes (e.g., “chair”, “car”), where shape plays a dominant role in visual discrimination. This demonstrates that structural priors learned from minimal representations translate effectively to tasks requiring precise spatial reasoning.
These results extend our findings beyond classification and detection, highlighting that line drawing pretraining induces representations that generalize across both sparse and dense vision tasks.
Figure 8: Different line drawing pretraining methods on ResNet18 generalizing to segmentation on ADE20K.
Benefits in Object Detection
Here, we also run similar experiments in Figure 3 on object detection task using PASCAL VOC dataset. Figure 9 illustrates that models pretrained on line drawings consistently outperform color-only models across all training data regimes. Notably, our Linedrawing→Photo (Ours) model trained with just 90% of the data achieves an mAP of 56.74, surpassing the Photo Only model trained with the full 100% dataset, which reaches 56.14. This highlights a key advantage of our approach: better performance with less data. The performance gap is particularly prominent in the low-data regime—for example, at 10% training data, our model achieves 22.16 mAP compared to 18.17 for the color-only baseline. These findings suggest that pretraining on structure-focused line drawings enables models to extract more efficient and transferable features for object detection, leading to significant gains in data efficiency and generalization.
@@
Compact and Transferable Representations
One of the core advantages of line-drawing pretraining is the emergence of compact and structurally focused internal representations. To investigate this, we first analyze the intrinsic dimensionality of neural activations in a ResNet-18 model trained under two regimes: color-only vs. line-drawing → color. We extract penultimate layer activations across STL10 test images and compute principal components (PCs) via PCA.
As shown in Figure 10, the line-drawing-pretrained model requires only 18 PCs to explain 90% of the variance, compared to 30 for the color-only model—suggesting that the model organizes information more economically and filters out redundant features. This compressed manifold structure aligns with principles of efficient coding and complements earlier findings on activation sparsity (Figure 6).
We further examine whether this structural compactness benefits model distillation. Using a standard teacher-student framework, we compare the performance of student models (MobileNetV1, ResNet-8, VGG-8) distilled from ResNet-18 teachers trained with either color-only or line-drawing → color supervision. As shown in Table 3, students distilled from line-pretrained teachers consistently outperform those trained from color-only teachers. For example, a MobileNetV1 distilled from a line-drawing-pretrained teacher (Pretrained with 70%-data to reduce the confounding factor of a stronger teacher) achieves 75.48% accuracy—exceeding the 73.83% achieved by the color-only teacher while the two teacher share almost the same object recognition performance.
These results suggest that line-drawing pretraining not only reduces the complexity of internal representations, but also produces knowledge that is easier to transfer, yielding more performant student models.
@@
[…]
Benefits of Unsupervised Line Drawing Generation
Training Strategy Color Image Accuracy
Color Training Only 68.79%
[…]
Learning to Draw $\rightarrow$ Color Training 73.30%
Table 4: Comparison of unsupervised Learning to Draw pretraining on COCO and then finetuning on STL10 color image classification accuracy.
Our unsupervised Learning to Draw approach—training models to generate line drawings from color images without paired supervision—significantly improves visual recognition. The results presented in Table 4 demonstrate the effectiveness of our proposed unsupervised Learning to Draw approach as a pretraining strategy for visual recognition. It achieves 73.30% accuracy on STL10, surpassing color-only training by 4.51 points and reaches almost the same level as the supervised line-drawing pretraining. These results validate our hypothesis: learning to synthesize line drawings guides models to extract structural, shape-centric representations, even without explicit supervision.
Conclusion
@@
Our findings show that line drawing pretraining induces compact, structure-aware representations that generalize across tasks and enable more effective distillation into lightweight models. This structure-first learning approach improves performance, enhances data efficiency, and yields representations that are easier to compress and transfer. More broadly, our results offer a computational lens on the longstanding observation that humans can easily understand line drawings—suggesting that prioritizing geometry over appearance may support more general and efficient visual reasoning, both in artificial and biological systems.

### v2

@@ 4.2 Data Efficiency
Our experimental results provide compelling evidence for the data efficiency advantages of line drawing curriculum learning over conventional color-only training approaches. As illustrated in Figure 3, the Line-Color method demonstrates superior learning efficiency across all training data regimes, achieving the baseline color accuracy (69%) with only 70% of the training data—representing a substantial 30% reduction in labeled data requirements. This efficiency gain manifests early in the training process, with the Line-Color approach consistently outperforming the color-only baseline by 3-5 percentage points at equivalent data volumes. Notably, the learning trajectory exhibits a steeper ascent during the initial training phases, suggesting that the line drawing curriculum provides a more structured optimization landscape that facilitates faster convergence to target performance levels.
The comparative analysis with style controls (Matisse-Color, Mondrian-Color, and Picasso-Color, Canny-Color) further validates the specificity of line drawing advantages. We specifically include Canny edge images as a naive control, as they represent a classic, low-level approach to structure extraction that lacks the semantic coherence of learned line drawings. While some of these alternative visual abstractions yield improvements over the color-only baseline, they are significantly less effective than line drawings in achieving data efficiency. The other methods require substantially more training data to reach baseline performance compared to the line drawing approach. This differential performance strongly supports our theoretical framework that line drawings tap into fundamental shape and topology processing mechanisms inherent to visual recognition systems. The reduced effectiveness of alternative abstractions demonstrates that the efficiency gains are not merely artifacts of stylistic variation, but rather reflect the unique capacity of line drawings to provide structured, shape-preserving visual representations that align with the computational principles underlying human visual processing.
Furthermore, these data efficiency benefits extend beyond convolutional architectures. As demonstrated with a Vision Transformer (ViT-Tiny) in Figure 4, the advantages of the structure-first curriculum are even more pronounced. The ViT model trained with the Line-Color curriculum achieves the final performance of the color-only baseline while using approximately 40% less training data, a greater reduction than observed with ResNet-18. This consistent outperformance across both CNN and transformer-based models underscores the general applicability of our approach, confirming that the induced structural priors provide a robust, cross-architecture benefit for efficient visual learning.
4.3 More Focused Attention
Figure 5: GradCAM visualizations comparing attention heat maps for models trained on color images vs trained with our curriculum.
Figure 6: Neuron tuning curves of base ResNet-18 vs ResNet-18 with our curriculum. (Normalized activations over 500 test images)
To understand the mechanisms driving these performance improvements, we analyzed the attention patterns of models trained under different learning curricula. Figure 5 presents GradCAM (jacobgil:pytorchcam) visualizations comparing models trained exclusively on color images versus those trained with our structure-first curriculum (STL10-Line followed by STL10 color).
Models initialized with structure-first learning on line drawings demonstrate markedly more focused attention patterns, with heat maps concentrated on the primary subject. For the cat image, the structure-first model concentrates on facial features and body contours, while the color-only model’s activation is more diffuse. Similarly, for the airplane, the structure-first model focuses on the aircraft’s structural elements, unlike the scattered activation of the color-only model. A more quantitative analysis of these GradCAM activations is included in the supplementary material.
4.4 Sparse Neuronal Tuning Curves
The neuron tuning curve presented in Figure 6 provides quantitative support for these qualitative observations. We evaluate the activation tuning curve for the last layer of the trained ResNet-18. Each neuron in Figure 6 represents one channel. We take the average firing activations from each channel and normalize them according to their maximum firing activation within each setting. The resulting activation tuning curve shows that ResNet-18 first trained on line drawings and then adapted to natural images (blue curve) exhibits a much sharper tuning curve in activation compared to the baseline counterpart (red curve). This pattern indicates that strong responses are concentrated among a small subset of neurons, while most units remain weakly or sparsely engaged. Such a rapid fall-off shows an increased sparsity and efficiency in neural representations.
Instead of distributing activation broadly, the network from the structure-first curriculum allocates capacity to the most informative neurons, silencing irrelevant or noisy features. The base model, by contrast, shows a more gradual decrease, activating a broader collection of neurons, including many that likely encode non-discriminative or redundant information.
This selective focus echoes principles of efficient coding: the model learns to emphasize features that matter for object recognition while suppressing spurious or noisy activations. This transition to more selective, quieter representations improves performance without any trade-off. The structure-first model achieves stronger accuracy on both minimal and natural images.
These results encapsulate the representational transformation induced by an initial phase of minimal, structural learning, which encourages the model to specialize its detectors for geometric cues.
4.5 Increased Shape Bias
@@
To further assess the inductive biases, we evaluated shape versus texture preferences using the model-vs-human benchmark (geirhos:2021). This benchmark presents cue-conflict images in which shape and texture cues are intentionally mismatched, allowing for a direct comparison of classification strategies between models and human observers.
Figure 7 displays the distribution of shape and texture decisions across object categories (these are the 7 categories in the cue-conflict dataset that coincide with STL10 classes). Models first trained on line drawings and subsequently adapted to color images demonstrate a pronounced shift toward shape-based decisions, as indicated by the clustering of yellow circles toward higher fractions of shape choices. In contrast, models trained only on natural images show a greater reliance on texture cues, with more blue circles distributed toward the texture-dominant end of the spectrum. This pattern suggests that an initial phase of learning on line drawings encourages the model to prioritize global structural features over local textural information when resolving ambiguous visual input.
These results indicate that our structure-first approach not only enhances classification performance and attention focus, but also aligns model decision-making more closely with human perceptual strategies. This increased shape bias likely contributes to the improved robustness and generalization observed in downstream tasks, as models become less reliant on superficial texture cues and more attuned to the underlying structure of visual objects.
4.6 Benefits in Semantic Segmentation
We further show that an initial phase of learning on line drawings enhances generalization to dense prediction tasks such as semantic segmentation. To evaluate this, we adapt ResNet-18 models—initialized either from color-only training or from our structure-first curriculum—on the ADE20K benchmark.
As shown in Figure 8, the model initialized with line drawings consistently outperforms the color-only baseline across multiple semantic categories. The improvement is especially pronounced for object-centric classes (e.g., “chair”, “car”), where shape plays a dominant role in visual discrimination. This demonstrates that structural priors learned from minimal representations translate effectively to tasks requiring precise spatial reasoning.
These results extend our findings beyond classification and detection, highlighting that our structure-first curriculum induces representations that generalize across both sparse and dense vision tasks.
Figure 8: Different learning curricula on ResNet-18 generalizing to segmentation on ADE20K.
4.7 Benefits in Object Detection
We also run experiments similar to those shown in Figure 3, but on the object detection task using the PASCAL VOC dataset. Figure 9 illustrates that models trained with our structure-first curriculum consistently outperform color-only models across all training data regimes. Notably, our Linedrawing→Photo (Ours) model trained with just 90% of the data achieves an mAP of 56.74, surpassing the Photo Only model trained with the full 100% dataset, which reaches 56.14. This highlights a key advantage of our approach: better performance with less data. The performance gap is particularly prominent in the low-data regime—for example, at 10% training data, our model achieves 22.16 mAP compared to 18.17 for the color-only baseline. These findings suggest that an initial learning phase on structure-focused line drawings enables models to extract more efficient and transferable features for object detection, leading to significant gains in data efficiency and generalization.
@@ 4.7 Benefits in Object Detection
4.8 Compact and Transferable Representations
One of the core advantages of a structure-first curriculum is the emergence of compact and structurally focused internal representations. To investigate this, we first analyze the intrinsic dimensionality of neural activations in a ResNet-18 model trained under two regimes: color-only vs. line-drawing → color. We extract penultimate layer activations across STL10 test images and compute principal components (PCs) via PCA.
[…]
As shown in Figure 10, the model from the structure-first curriculum requires only 18 PCs to explain 90% of the variance, compared to 30 for the color-only model—suggesting that the model organizes information more economically and filters out redundant features. This compressed manifold structure aligns with principles of efficient coding and complements earlier findings on activation sparsity (Figure 6).
@@ 4.8 Compact and Transferable Representations
Table 4: This table compares knowledge distillation from a line-drawing → color teacher (2) and a standard color teacher (3). For a fair comparison, their accuracies are matched by training (2) on 70% of the data while (3) uses 100%. The line-drawing teacher (2) yields superior student models. For reference, we also show the performance of an unmatched, higher-performing line-drawing teacher (1), which was trained on 100% of the data.
We further examine whether this structural compactness benefits model distillation. Using a standard teacher-student framework, we compare the performance of student models (MobileNetV1, ResNet-8, VGG-8) distilled from ResNet-18 teachers trained with either color-only or line-drawing $\rightarrow$ color supervision. As shown in Table 4, students distilled from teachers developed with the structure-first curriculum consistently outperform those from color-only teachers. For example, a MobileNetV1 distilled from a teacher from the structure-first curriculum (trained with 70%-data to reduce the confounding factor of a stronger teacher) achieves 75.48% accuracy—exceeding the 73.83% achieved by the color-only teacher while the two teachers share almost the same object recognition performance.
These results suggest that the structure-first curriculum not only reduces the complexity of internal representations, but also produces knowledge that is easier to transfer, yielding more performant student models.
5 Conclusion
@@
Our findings show that a structure-first curriculum using line drawings induces compact, structure-aware representations that generalize across tasks and enable more effective distillation into lightweight models. This structure-first learning approach improves performance, enhances data efficiency, and yields representations that are easier to compress and transfer. More broadly, our results offer a computational lens on the longstanding observation that humans can easily understand line drawings—suggesting that prioritizing geometry over appearance may support more general and efficient visual reasoning, both in artificial and biological systems.
Appendix A Appendix
A.1 Supervised Learning Statistical Analysis
Data (%) Color Only Line→Color Matisse→Color Mondrian→Color Picasso→Color
5 $30.62\pm 0.77$% $2.95\pm 0.5933$$3.79\pm 0.92$% $30.49\pm 1.87$% $32.64\pm 0.68$%
10 $36.25\pm 1.66$% $37.68\pm 0.30$% $38.40\pm 2.34$% $38.76\pm 0.94$% $39.25\pm 1.56$%
20 $45.78\pm 3.22$% $51.12\pm 2.96$% $43.50\pm 1.73$% $47.64\pm 2.48$% $45.80\pm 3.15$%
30 $49.51\pm 4.10$% $57.48\pm 1.08$% $52.01\pm 3.67$% $50.70\pm 1.42$% $52.04\pm 2.19$%
40 $56.53\pm 0.75$% $60.33\pm 1.58$% $58.10\pm 0.84$% $60.59\pm 2.73$% $58.24\pm 1.29$%
50 $58.47\pm 1.27$% $62.89\pm 1.60$% $61.76\pm 2.05$% $62.24\pm 0.76$% $62.64\pm 2.81$%
60 $60.76\pm 1.87$% $67.75\pm 1.15$% $3.04\pm 1.3866$$6.12\pm 2.94$% $63.95\pm 0.89$%
70 $66.49\pm 0.71$% $69.15\pm 4.03$% $4.78\pm 2.6766$$6.99\pm 1.03$% $65.50\pm 1.85$%
80 $66.84\pm 2.09$% $9.48\pm 2.1366$$6.94\pm 0.95$% $67.55\pm 1.74$% $67.24\pm 2.36$%
90 $67.21\pm 2.34$% $71.51\pm 0.46$% $19\pm 3.0868.68$$68\pm 0.62$% $68.00\pm 1.47$%
100 $68.79\pm 0.91$% $73.58\pm 0.42$% $68.77\pm 1.53$% $71.40\pm 0.56$% $69.03\pm 0.77$%
Table 5: Aggregated results with std dev across all seeds showing performance comparison between Color Only training and different initial learning modalities.
To rigorously evaluate the statistical significance of our findings, we conducted Wilcoxon signed-rank tests comparing each initial learning modality against the Color Only baseline. The Wilcoxon signed-rank test is a non-parametric statistical test that assesses whether there is a significant difference between paired observations without assuming a normal distribution, making it particularly appropriate for our experimental setup where we have matched pairs of performance measurements across different data percentages.
For each comparison, we tested the null hypothesis $H_{0}$: median difference = 0 against the alternative hypothesis $H_{1}$: method performance >> Color Only baseline (one-tailed test). The test was applied to 11 paired observations corresponding to the performance differences at each data percentage level (5%, 10%, 20%, …, 100%).
Statistical Results: Line→Color vs Color Only shows a median improvement of +4.33 percentage points (W-statistic: 66, p-value: 0.002, highly significant), with Line→Color outperforming Color Only in all 11 data conditions (11/11 wins). Other stylistic augmentations showed: Mondrian→Color +1.47 percentage points (p = 0.037, significant), Picasso→Color +0.24 percentage points (p = 0.195, not significant), and Matisse→Color -0.02 percentage points (p = 0.500, not significant).
Implications: The statistical analysis provides strong evidence for several key conclusions. The Line→Color approach demonstrates highly significant improvement over the Color Only baseline (p = 0.002), with consistent gains across all tested data percentages. This statistical significance, combined with the perfect win rate (11/11), indicates that the performance benefit is both reliable and substantial. Critically, our results show that Line→Color’s benefits cannot be attributed merely to stylistic augmentation effects. While Mondrian→Color shows modest improvement (+1.47%, p = 0.037), both Matisse→Color and Picasso→Color fail to achieve statistical significance. The substantial difference between Line→Color’s improvement (+4.33%) and that of other stylistic methods suggests that line drawings possess unique structural properties that facilitate more effective learning. The consistent performance gains across all data percentages (5% to 100%) demonstrate that an initial phase of learning on line drawings provides benefits in both low-data and high-data scenarios. The median improvement of 4.33 percentage points represents a practically meaningful enhancement in classification performance that could translate to significant real-world benefits. These statistical findings support our central hypothesis that line drawings provide structurally advantageous representations for subsequent learning phases, offering benefits that extend beyond those achievable through general stylistic augmentation approaches.
A.2 Computing Infrastructure
Experiments are run on a computing cluster equipped with dual NVIDIA L40 GPUs and AMD EPYC 7443 24-Core Processors running at 2.84GHz. Each experimental job utilized 8 CPU cores and 32GB of system RAM. The computing environment uses Springdale Open Enterprise Linux 8.6 (Modena) operating system on x86-64 architecture.
A.3 Datasets Preparation
To generate the line drawings for our initial structural learning phase, the pre-existing “Anime style” model from https://github.com/carolineec/informative-drawings (chan:2022) was used to create line drawing versions of STL10 (coates:2011stl10) and ImageNet-1K (russakovsky2015imagenet).
For the other augmentations, code from https://github.com/xunhuang1995/AdaIN-style (huang:2017) was used to generate Matisse, Mondrian, and Picasso style STL10 images with degree of stylization $\alpha=0.6$.
The base ADE20K (zhou2017ade20k) and Pascal VOC (pascalvoc2010) are also used downloaded for evaluation on segmentation and detection.
A.4 Supervised Learning Experiments
We conduct experiments comparing two training approaches: a baseline Color-Only Training (direct training on STL10 color images) and our Line→Color curriculum (an initial phase of learning on line drawings followed by a second phase on color images). As a control, we replicated this same curriculum with 3 different arbitrary augmentations (Matisse, Mondrian, Picasso) to show that performance benefits came from the intrinsic properties of line drawings and not just the effects of using an alternative visual modality for an initial learning stage. The dataset is STL10 with 10-class image classification, using 80% train / 20% validation split with stratified sampling. We test data percentages from 5% to 100% in increments of 5% or 10%.
The model architecture uses ResNet-18 trained from scratch, modified with a final linear layer (512 → 10) for 10-class classification, totaling approximately 11.2M parameters.
During hyperparameter development, we explored learning rates from 0.0001 to 0.01, batch sizes from 4 to 64, early stopping patience from 5 to 15 epochs, and weight decay from 1e-5 to 1e-3. Final hyperparameters use adaptive learning rates and batch sizes based on dataset size: 0.001/4 for small datasets ($\leq$300 samples), 0.003/8 for medium datasets ($\leq$1000 samples), 0.005/16 for larger datasets ($\leq$2000 samples), and 0.01/32 for full datasets (>>2000 samples).
Training configuration uses SGD optimizer with momentum 0.9, ReduceLROnPlateau scheduler with patience 3, early stopping with patience 10 epochs, and maximum 100 epochs. The primary evaluation metric is test accuracy on held-out test sets.
All experiment runs use system-generated randomness with no fixed seeds, reflecting natural variation and real-world training variability. Experimental runs use 5 runs per configuration with different random seeds. The training code is available in the code appendix in Training/train_resnet18.py.
We also conducted experiments using Vision Transformer Tiny (ViT-Tiny) architecture for comparison with the ResNet-18 baseline. The model uses the vit_tiny_patch16_224 configuration from the timm library, trained from scratch on STL10 with 10-class classification, containing approximately 5.7M parameters with patch size 16x16 and input resolution 224x224. The architecture includes 12 transformer layers, 3 attention heads, hidden dimension 192, learnable position embeddings, and a classification head ($192\to 10$) for 10-class classification.
The training configuration was optimized specifically for small datasets based on recommendations from (steiner2022trainvitdataaugmentation). We used AdamW optimizer with learning rate 0.0005, weight decay 0.01, and betas (0.9, 0.999), combined with cosine annealing learning rate schedule and 10% warmup epochs. The batch size was set to 32 (reduced from standard 64 for better gradient estimates on small datasets) with light augmentation strategy including RandomResizedCrop (scale 0.85-1.0), RandomHorizontalFlip (p=0.5), and mild ColorJitter. Regularization included conservative mixup ($\alpha=0.1$, applied 30% of the time), label smoothing (0.05), and gradient clipping (max_norm=0.5). Training used maximum 150 epochs with early stopping patience of 20 epochs.
The configuration specifically addresses challenges of training ViTs on small datasets through reduced regularization, conservative augmentation, smaller batch sizes, and extended patience periods. The curriculum learning approach follows the same Line→Color transfer strategy as the ResNet-18 experiments, with Xavier uniform initialization for linear layers and standard initialization for LayerNorm components following best practices for training ViTs from scratch. Training code is available in the code appendix in Training/train_tiny_vit.py
A.5 GradCAM Distribution Analysis
We analyze the spatial distribution of model attention by counting distinct high-activation regions and average size of said regions in GradCAM heatmaps. This quantifies whether models focus on concentrated areas or distribute attention across multiple regions.
For each GradCAM heatmap, we compute a dynamic threshold as the 85th percentile of heatmap values. We then create a binary mask where pixels above the threshold are marked as attention regions. To remove noise, we filter out regions smaller than 10 pixels and count distinct connected components using 8-connectivity labeling.
Figure 11: Example of counting number of high activation regions from GradCam. The GradCAM heatmap for base ResNet-18 has 5 distinct activation regions on this image while the ResNet-18 from the structure-first curriculum has one larger focused activation region.
The attention scatter metric is defined as the number of distinct attention regions per image. We visualize the distribution of region counts as a histogram to show the frequency of different attention scatter patterns.
For mean region size, we calculate the area of each connected component in pixels and compute the average region size per image. This metric indicates whether attention regions tend to be large and concentrated or small and scattered. Quantitative analysis confirms that the structure-first curriculum leads to a larger number of images with a single, compact attention region, whereas color-only training results in more fragmented attention (Figure 12). The GradCam analysis is available in the code appendix in Evaluation/gradcam_activations.py.
A.6 Neuron Tuning Curve Analysis
The neuron tuning curve analysis examines how different initial learning modalities affect the activation patterns of individual channels in the final convolutional layer of ResNet-18 models. In this analysis, each ”neuron” corresponds to one of the 512 channels in the final convolutional layer (layer4) of ResNet-18. The activation of each neuron is computed by averaging the spatial dimensions (height and width) of the corresponding feature map.
The analysis processes a dataset of test images (up to 500 images) through both models: ResNet-18 trained on color images only (Color Only), and ResNet-18 first trained on line drawings followed by a second phase of training on color images (Line→Color). For each image, we extract the 512-dimensional neuron activation vector using forward hooks registered on the layer4 module.
For each model, we compute the mean activation across all processed images for each of the 512 neurons. The neurons are then sorted by their mean activation values in descending order, creating a tuning curve that shows the distribution of neuron response strengths. The curves are normalized by dividing by the maximum activation value to facilitate comparison between models.
The resulting tuning curves reveal the degree of specialization in the learned representations. A steeper, more concentrated curve indicates that fewer neurons are highly active, suggesting more specialized, sparse representations. By comparing the curves between Color Only and Line→Color models, we can assess whether an initial phase of learning on line drawings leads to more efficient, specialized neuronal representations. The code for this is available in the code appendix in Evaluation/neuron_tuning_curves.py.
A.7 Generalizing to Segmentation
To evaluate how well ResNet-18 models initialized with different visual modalities generalize to semantic segmentation, we conducted experiments on the ADE20K dataset using a standard encoder-decoder architecture. We compared two ResNet-18 encoders developed through different curricula: one trained on color images and another trained sequentially on line drawings followed by color images (Line→Color). We also include another ResNet-18 that is simultaneously trained on line and color images to investigate whether training order affects generalization ability.
We modified the semantic segmentation framework from https://github.com/CSAILVision/semantic-segmentation-pytorch (zhou2018semantic) to support our custom encoders. The training configurations closely match the original framework’s settings to ensure a fair comparison. We employed a Pyramid Pooling Module (PPM) decoder architecture with ResNet-18 as the encoder backbone.
The training setup was identical for all models to ensure fair comparison. We trained for 20 epochs with a batch size of 2 per GPU using SGD optimization. The learning rate was set to 0.02 for both encoder and decoder, with a polynomial decay schedule ($lr_{current}=lr_{initial}\times(1-\frac{iter}{max\_iter})^{0.9}$). Both the encoder and decoder were trained end-to-end, allowing the learned representations to be adapted for the segmentation task.
The models were trained on ADE20K with 150 semantic classes, using images resized to multiple scales $(300,375,450,525,600)$ with a maximum size of 1000 pixels. We employed standard data augmentation including random horizontal flipping and used the negative log-likelihood loss with ignored index -1 for handling invalid labels.
Both models were evaluated using the same checkpoint at epoch 20. The evaluation metrics include pixel accuracy and mean intersection-over-union (mIoU) on the ADE20K validation set. This experimental setup allows us to directly compare the representational quality of the two learning curricula by measuring their performance on a downstream semantic segmentation task. The code for this is available in the code appendix in Training/Segmentation.
A.8 Object Detection in PASCAL VOC
We utilize the PASCAL VOC dataset to evaluate data efficiency on object detection. For our curriculum, we first train the model on line drawings for 200 epochs and then train on color images for another 200 epochs. The base model is trained on photographic images alone for 200 epochs. We follow the training setup from ultralytics reis2023real and use the YOLO-v8-n architecture. The code to run experiments in this section is located under Training/ObjectDetection.
A.9 Knowledge Distillation Experiments
We utilize the STL10 classification task for knowledge distillation. The codebase is included in the supplementary folder Training/KnowledgeDistillation as well as the experiment setup. The experiments used three types of teachers: (1) The best performing model developed through our Line→Color curriculum; (2) The model developed with the Line→Color curriculum, where both learning stages are conducted using only 70% of the data. With less data, the model performance drops but still matches the baseline training with raw color images on 100% data. Using such a teacher ensures a fair comparison between the model from our structure-first curriculum and the baseline model. The experimental results in Table 4 in the main manuscript suggest that the line drawings produce a better representation that is superior when used for knowledge distillation. (3) The baseline model trained on 100% of the training data.
A.10 Code
Source code for this project can be found at the following anonymized repository: https://github.com/GeorgeLiu59/Learning-More-Seeing-Less.